### v8

@@ Abstract
Due to the dynamic nature of the Semantic Web, version control is necessary to manage changes in widely used ontologies. Despite the long-standing recognition of ontology versioning (OV) as a crucial component of efficient ontology management, many approaches treat OV as similar to ontology matching (OM) and directly reuse OM systems for OV tasks. In this study, we systematically analyse similarities and differences between OM and OV and formalise the OM4OV pipeline to offer more advanced OV support. The pipeline is implemented and evaluated in the state-of-the-art OM system Agent-OM. The experimental results indicate that OM systems can be reused for OV tasks, but without necessary extensions, the current OM4OV pipeline can produce skewed measurements, poor performance in detecting update entities, and limited explainability of false mappings. To tackle these issues, we propose an optimisation method called the cross-reference (CR) mechanism, which builds on existing OM alignments to reduce the number of matching candidates and to improve overall OV performance.
@@ 1 Introduction
Ontologies serve as the backbone of the Semantic Web, providing formal concept descriptions of shared concepts across various applications [14]. An ontology is not static, and the need for version control arises. While web data is dynamic, any ontology used must undergo periodic revisions to keep pace with the growth in domain knowledge, modifications to the application adaptation, or corrections to the shared conceptualisation [19]. For example, it is unrealistic to expect ontologies created in the 1990s to contain concepts such as “touchscreen”, “fingerprint sensor”, or “WiFi antenna” [15]. These modifications may cause undesirable deficiencies in artifacts that conform to or reuse the ontology being changed, leading to severe non-compliance and incompatibility issues in downstream tasks.
Ontology versioning (OV), also known as ontology evolution, aims to distinguish and recognise changes between different ontology versions. By doing so, data that conforms to the changed ontology, ontologies that reuse the changed ontology, or software that uses the changed ontology can apply the correct changes correspondingly [19]. Although OV plays an important role in ontology management, this area remains underexplored. Many studies [27, 26, 28, 7, 12, 34, 21, 22, 23, 30] treat OV as similar to ontology matching (OM) and assume OM systems can be reused for OV tasks. OM is a well-studied problem, so leveraging this deep research legacy for OV is attractive (we call this approach “OM4OV”).
In this paper, we assume the Web Ontology Language (OWL) [5] is used to present an ontology. We use “concept” and “entity” to refer to OWL classes or properties. We use “subsumption” to mean “subclass of” and “equivalence” to mean “equivalent” in OWL as formal concept inter-relations. We analyse the complementary relationship between the OM and OV tasks and provide a task formulation for the OM4OV pipeline. We implement the pipeline in a modern OM system to validate its effectiveness in unifying the OM and OV tasks. We investigate the links between these two tasks in practice and explore optimisations for OM4OV. To the best of our knowledge, this work is the first to systematically analyse the OM4OV pipeline and provide theoretical foundations and empirical evidence for its effectiveness and pitfalls. Specifically, our key contributions include:
  • We analyse the similarities and differences between OM and OV. In the OM4OV pipeline, we describe four disjoint subset alignments produced in the OV process: remain, update, add, and delete, each corresponding to a category of match or non-match. We provide formal definitions and formulations.
  • We implement the OM4OV pipeline in the state-of-the-art OM system Agent-OM [32] and its extension Agent-OV and evaluate the performance on the testbed derived from the Ontology Alignment Evaluation Initiative (OAEI) [8] datasets. We find that OM systems can be reused for OV tasks, but extensions are necessary to improve overall OV performance and its meaningful quality measurement.
@@ 2 Related Work
Maintaining version information in the ontology can be labour-intensive and error-prone. Either extending the current schema or using change logs requires consistent updating over time. In most cases, this process is hand-crafted by the ontology engineer or requires human intervention (e.g. pre-defining a schema or creating a template for the change log). A manual process is more likely to make mistakes and fail to propagate changes to dependent artefacts. Also, in the real world, there is no guarantee that the ontology itself contains the complete version information or that it has a separate change log. In such cases, current approaches have limited capability to detect incorrect version information or to identify missing version information. These approaches rely on the version information contained in or attached to the ontology by manual methods. If such information is missing or incorrect, automated versioning of ontology concepts is necessary.
@@
Reusing OM systems for OV tasks (OM4OV) is a lightweight and fully automatic approach. OM4OV is compatible with ontologies using different recording methods, including URIs or additional versioning triples, as well as ontologies that lack or do not include version information. The PromptDiff suite [27, 26, 28] is the canonical work that integrates different heuristic matchers and applies a matching-like comparison for OV. The concept of “matching” is later used in the H-Change methodology [7] to assess the assimilation of new concepts into the old ontology. While studies in [12, 34] highlight that changes in ontologies can lead to corresponding changes in ontology mapping sets, the authors in [21, 22, 23] state that updating matching alignments is the key to tracking ontology version changes. A recent work in [30] develops an ontology evolution maintenance framework that combines a set of algorithms to revalidate ontology alignments. However, existing naive OM4OV approaches directly use OM systems for OV and focus on developing new algorithms and architectures for ontology version control, and less attention has been paid to understanding the underlying differences between OM and OV. OM is not exactly the same as OV. OM determines matched entities between two different ontologies, while OV distinguishes add, delete, remain, and update entities over different versions of one ontology. The OM input consists of two distinct ontologies, whereas the OV input is expected to consist of two versions of a single ontology. The output of OM is a set of entity mappings, whereas the output of OV comprises four sets, each defined by the nature of the change made. These differences may affect the adaptability of reusing OM systems for OV tasks.
3 OM4OV Task Formulation
In this section, we draw concepts in OM to formalise OV tasks.
@@ 3 OM4OV Task Formulation
OM Preliminary: Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), the OM task is to find an alignment $A$ (i.e. a set of mappings) with respect to a given similarity threshold $s\in[0,1]$, defined as [10]: $A=\{(e1,e2,r,c)|e1\in O_{s},e2\in O_{t},s\leq c<1\}$, where $e1$ and $e2$ are ontology entities in $O_{s}$ and $O_{t}$ respectively, $r$ states the relation between $e1$ and $e2$ which can be equivalence ($\equiv$) or subsumption ($\subseteq$), and $c\in[0,1]$ is the level of confidence for the match $(e1,r,e2)$.
@@
Given an old version of an ontology ($O$) and a new version of the same ontology ($O^{\prime}$), an OV task can be considered as finding an alignment (A) that contains a collection of mappings over the two different versions. As illustrated in Figure 1, while an alignment for OM only considers matched entities, OV tracks both matched and non-matched entities. Furthermore, matched entities are composed of two subsets remain and update, while non-matched entities are composed of add and delete entities.
Figure 1: Overview of the OM4OV pipeline.
@@
An OV task can be formalised as finding two variants of an alignment $A_{match}$ and $A_{non-match}$ between ontology versions $O$ and $O^{\prime}$ with respect to a given similarity threshold $s\in[0,1]$. OV focuses only on the equivalence relation, and subsumption mappings are beyond the scope of versioning. We assume the relation $r$ is an equivalence and do not refer to the relation henceforward. Therefore, the OV task is defined as:
$A_{match}$ $=\{(e1,e2,c)|e1\in O,e2\in O^{\prime},s\leq c\leq 1\}$ (1)
$A_{non-match}$ $=\{(e1,e2,c)|e1\in O,e2\in O^{\prime},0\leq c<s\}$
@@
The remain and update entities are actually matched entities between different ontology versions. The remain entities can be considered to be the exact (trivial) equivalent string matches between different ontology versions, assuming that the naming convention does not change between different ontology versions. If the naming convention changes, we will classify ontology entities that have changed names according to the systematic convention as update entities, by modelling their confidence $c$ as a value greater than $s$ but less than 1. The update entities are, in general, the closest (non-trivial) near-equivalent matches between different ontology versions. Therefore, $A_{remain}(A\odot)$ and $A_{update}(A\otimes)$ are defined as:
$A_{match}=A\odot\cup A\otimes$ (2)
$A\odot=A_{match}(c=1)=\{(e1,e2,c)|e1\in O,e2\in O^{\prime},c=1\}$
$A\otimes=A_{match}(s\leq c<1)=\{(e1,e2,c)|e1\in O,e2\in O^{\prime},s\leq c<1\}$
Definition 3.4.
The add and delete entities are actually non-matched entities between different ontology versions. The add entities are the non-matched terms in $O^{\prime}$, which have no matches in $O$. Similarly, we can interpret the delete entities as non-matched entities in $O$. Without loss of generality, we consider the set of entities in the add and delete categories to be alignments by equivalently interpreting them. Therefore, $A_{add}(A\oplus)$ and $A_{delete}(A\ominus)$ are defined as:
$A_{non-match}=A\oplus\cup A\ominus$ (3)
$A\oplus=\{e2\in O^{\prime}\cap A_{non-match}\}=\{e2\in O^{\prime}\setminus A_{match}\}$
$A\ominus=\{e1\in O\phantom{{}^{\prime}}\cap A_{non-match}\}=\{e1\in O\phantom{{}^{\prime}}\setminus A_{match}\}$
@@ 4.1 Dataset Construction
There is a dearth of benchmark datasets for evaluating OV. The Ontology Alignment Evaluation Initiative (OAEI) contains several datasets related to OM tasks, but none are specifically designed for OV tasks. We propose an approach to constructing synthetic OV datasets from OM datasets. Figure 2 illustrates the generation of OM4OV datasets.
Figure 2: Generation of OM4OV datasets.
@@
The original OAEI datasets for OM provide two ontologies, the source $O_{s}$ and the target $O_{t}$. We choose one as the ontology for versioning and duplicate it into $O$ and $O^{\prime}$. Another ontology will be used later to provide references for update entities. We duplicate this ontology into $O_{r}$.
There are four possible entity changes in OV tasks: remain, update, add, and delete. We retrieve all ontology entities from $O$ or $O^{\prime}$ (since they are identical at this stage). Each ontology entity will be randomly assigned to one of these categories, and the four sets are pairwise disjoint by construction. In reality, there are limited entities that have replacements. To ensure that the percentage of update entities can reach 25% of the distribution, we prioritise the assignment of update entities and assign the other three categories until the maximum replacement or 25% ontology size is reached.
For each ontology entity $e$ assigned to a category, the corresponding dataset operations are described as follows. While there is no operation for remain entities, entities assigned to add and delete are required to delete related triples in $O$ or $O^{\prime}$. The update entities need to replace all related triples in $O^{\prime}$. We generate four corresponding versioning references based on the entity assignments. These files provide the ground truth for evaluation.
To ensure the replaced entity is reasonable and reflects real-world facts, we use equivalent classes or properties specified in reference.xml. For example, replace cmt:ProgramCommitteeChair with its equivalent class confof:Chair_PC. This operation is cascaded, meaning that we update not only the entity name but also its related triples, such as changing the semantic information from (cmt:ProgramCommitteeChair, rdfs:subClassOf, cmt:ProgramCommitteeMember) to (confof:Chair_PC, rdfs:subClassOf, confof:Person). While this procedure can be unlimited, we restrict it to extract only the triples directly related to the entity (i.e. a 1-hop subgraph) in the ontology $O_{r}$. For entities whose names are unique identifiers or codes (and not textually meaningful names), we retrieve their meaningful descriptions from annotation properties. For example, we extract an additional annotation triple (mouse:MA_0000270, rdfs:label, “eyelid tarsus”) for the entity mouse:MA_0000270. After extraction, we also need to update the URI from “confof” to “cmt” where the new entity URI will be cmt:Chair_PC and the replacement triple will be (cmt:Chair_PC, rdfs:subClassOf, cmt:Person). Finally, we add these new triples to the existing ontology $O^{\prime}$. Table 1 summarises the detailed steps of the dataset operation for update entities.
@@
Unlike the original OAEI datasets used for OM, randomness ensures that the synthetic OAEI datasets for OV are different each time they are constructed. This suits the dynamic nature of OV, where the changes vary between different versions. For this reason, we consider the new OAEI datasets for OV more like a testbed, as they can simulate a variety of situations for OV tasks. For reproducibility, we use a fixed seed to control the randomness. Table 2 lists the detailed information of the selected OAEI track used for the OV testbed. The current version of the OV testbed contains a total of 11 distinct ontologies from three different OAEI tracks. The anatomy track contains two large ontologies, while the MSE track has two medium ontologies, and the conference track has 7 small ontologies. We exclude EMMO [9] in the MSE track because it uses a different naming convention than MaterialInformation [3] and MatOnto [18]. Figure 3 shows the number of entities in each track. The fixed seed for the OV testbed is set to 42.
@@
Track Domain # of Ontologies Anatomy Human and Mouse Anatomy 2 Conference Research Conference 7 MSE Materials Science & Engineering 2
@@
4.2 System Implementation
We choose the matching systems Agent-OM and its extended version Agent-OV for the experiment. Agent-OM is one of the flagship OM systems powered by large language models (LLMs) and multi-agent architectures. Its foundation framework is designed for traditional OM tasks. We extend Agent-OM with the OM4OV pipeline so that it can be used for OV tasks. Inherited from Agent-OM, Agent-OV supports a wide range of LLMs, including commercial API-accessed LLMs OpenAI GPT [29] and Anthropic Claude [2], as well as open-source LLMs Meta Llama [24], Alibaba Qwen [1], Google Gemma [11], and ChatGLM [36]. Similarly to Agent-OM, similarity_thresholdsimilarity\_threshold and $top@k$ are the adjustable key hyperparameters for Agent-OV.

### v9

@@ Abstract
Due to the dynamic nature of the Semantic Web, version control is necessary to manage changes in widely used ontologies. Despite the long-standing recognition of ontology versioning (OV) as a crucial component of efficient ontology management, many approaches treat OV as similar to ontology matching (OM) and directly reuse OM systems for OV tasks. In this study, we systematically analyse similarities and differences between OM and OV and formalise an OM4OV pipeline to offer more advanced OV support. The pipeline is implemented and evaluated in the state-of-the-art OM system Agent-OM. The experimental results indicate that OM systems can be effectively reused for OV tasks, but without necessary extensions, can produce skewed measurements, poor performance in detecting update entities, and limited explanation of false mappings. To tackle these issues, we propose an optimisation method called the cross-reference (CR) mechanism, which builds on existing OM alignments to reduce the number of matching candidates and to improve overall OV performance.
@@ 1 Introduction
Ontologies serve as the backbone of the Semantic Web, providing formal concept descriptions of shared concepts across various applications [14]. An ontology is not static, and the need for version control arises. While web data is dynamic, any ontology in use must undergo periodic revisions to keep pace with the growth in domain knowledge, modifications to the application adaptation, or corrections to the shared conceptualisation [19]. For example, it is unrealistic to expect ontologies created in the 1990s to contain concepts such as “touchscreen”, “fingerprint sensor”, or “WiFi antenna” [15]. These modifications may cause undesirable deficiencies in artifacts that conform to or that reuse the ontology being changed, leading to severe non-compliance and incompatibility issues in downstream tasks.
Ontology versioning (OV), also known as ontology evolution, aims to distinguish and recognise changes between different ontology versions. By doing so, data that conforms to the ontology, ontologies that reuse the ontology, or software that uses the ontology can apply the correct changes correspondingly [19]. Although OV plays an important role in ontology management, this area remains underexplored. Many studies [26, 25, 27, 7, 12, 34, 21, 22, 23, 29] treat OV as similar to ontology matching (OM) and assume OM systems can be reused for OV tasks. OM is a well-studied problem, so leveraging this deep research legacy for OV is attractive. We call this approach “OM4OV”.
We analyse the complementary relationship between OM and OV tasks and provide a task formulation for the OM4OV pipeline. We implement the pipeline in a modern OM system to validate its effectiveness in unifying the OM and OV tasks. We investigate the links between these two tasks in practice and explore optimisations for OM4OV. To the best of our knowledge, this work is the first to systematically analyse the OM4OV pipeline and to provide theoretical foundations and empirical evidence for its effectiveness and pitfalls. Specifically, our key contributions include:
  • We analyse the similarities and differences between OM and OV. In the OM4OV pipeline, we describe four disjoint subset alignments produced in the OV process: remain, update, add, and delete, each corresponding to a subset of match or non-match in OM. We provide formal definitions and formulations.
  • We implement an OM4OV pipeline in the state-of-the-art OM system Agent-OM [32] and its extension Agent-OV and evaluate the performance on the testbed derived from the Ontology Alignment Evaluation Initiative (OAEI) [8] datasets. We find that OM systems can be reused for OV tasks, but extensions are necessary to improve both overall OV performance and its meaningful quality measurement.
@@
In this paper, we assume the OWL 2 Web Ontology Language (OWL 2) [5] is used to present an ontology. We use both concept and entity to refer to OWL classes or properties. We use subsumption to mean rdfs:subClassOf or rdfs:subPropertyOf and equivalence to mean owl:equivalentClass or owl:equivalentProperty as formal class and property inter-relations.
@@ 2 Related Work
Version control is recognised as a vital element in ontology management. Different versions of an ontology need to support interoperability so that changes do not impede the effective and sustainable use of the ontology.
@@
Maintaining version information in the ontology can be labour-intensive and error-prone. Either extending the current standard 3-tuple schema or using change logs requires consistent updating over time. In most cases, this process is hand-crafted by the ontology engineer or requires human intervention (e.g. pre-defining a schema or creating a template for the change log). A manual process is more likely to make mistakes and fail to propagate changes to dependent artefacts. Also, in the real world, there is no guarantee that the ontology itself contains the complete version information or that it has a separate change log. In such cases, current approaches have limited capability to detect incorrect or missing version information. These approaches rely on the version information contained in or attached to the ontology by manual methods. If such information is missing or incorrect, automated versioning of ontology concepts is necessary.
Reusing OM systems for OV tasks (OM4OV) is a lightweight and fully automatic approach to detecting changes in ontologies. OM4OV is compatible with ontologies using different recording methods, including URIs or additional versioning triples, as well as ontologies that lack internal version information. The PromptDiff suite [26, 25, 27] is the canonical work that integrates different heuristic matchers and applies a matching-like comparison for OV. The concept of “matching” is later used in the H-Change methodology [7] to assess the assimilation of new concepts into the old ontology. Several studies [12, 34] note that updates to an ontology will generally create a requirement to update its existing alignments to external ontologies. Methods [21, 22, 23] are proposed to identify changes that affect alignments, and in some cases the framework is proposed to automate alignment updates [29], but these works do not address the discovery of alignments from an ontology to its own updated version.
@@
While existing OM4OV approaches directly use OM systems for OV and focus on developing new algorithms and architectures for ontology version control, paying less attention to understanding the underlying differences between OM and OV. OM is not exactly the same as OV. OM determines matched entities between two different ontologies, while OV can distinguish add, delete, remain, and update changes over different versions of one ontology. OM input consists of two distinct ontologies, whereas OV input is expected to consist of two versions of a single ontology. The output of OM is a set of entity mappings, whereas the output of OV comprises four sets, each defined by the nature of the change made. These differences may affect the adaptability of reusing OM systems for OV tasks.
3 Task Formulation
Now we draw on concepts in OM to formalise OV tasks. Given an ontology $O$ we say that some entity $e\in O$ when $e$ is of interest to ontology matching or ontology versioning, typically representing a concept of the domain of interest being modelled by $O$. We consider an entity to be the full IRI of the reference in an ontology, or its equivalent abbreviation via a prefix, for example cmt:ProgramCommitteeChair. We consider an entity name to be the substring of an entity that provides a meaningful description of the entity (e.g. ProgramCommitteeChair), excluding any entity prefix or expanded URI (e.g. not cmt:ProgramCommitteeChair).
3.1 OM Preliminary
@@
Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), the OM task is to find an alignment $A$ (i.e. a set of mappings) with respect to a given similarity threshold $s\in[0,1]$, defined as [10]: $A=\{(e_{1},e_{2},r,c)\mid e_{1}\in O_{s},e_{2}\in O_{t},s\leq c<1\}$, where $e_{1}$ and $e_{2}$ are ontology entities in $O_{s}$ and $O_{t}$ respectively, $r$ is the relation between $e_{1}$ and $e_{2}$ which is equivalence ($\equiv$) or subsumption ($\preceq$), and $c\in[0,1]$ is the level of confidence for the match $(e_{1},e_{2},r)$.
3.2 OM4OV Pipeline
Given an old version of an ontology ($O$) and a new version of the same ontology ($O^{\prime}$), the OV task can be considered as finding an alignment (A) from the old ontology to the new one. As illustrated in Figure 1, while an alignment for OM only mentions matched entities, OV needs to recognise both matched and non-matched entities. Furthermore, OV matched entities are comprised of two subsets: remain and update, while non-matched entities are comprised of add and delete entities. While some of these sets are comprised of entities, and others of mappings, when convenient henceforward, we will loosely refer to all of these as either alignments or sets of entities interchangeably.
Figure 1: Overview of the OM4OV pipeline. The sets are comprised of entities, with the matched intersection containing entities that are considered equivalent by an ontology alignment.
An OV task can be formalised as finding an alignment and a non-alignment, $A_{match}$ and $A_{non-match}$ respectively, between different ontology versions $O$ and $O^{\prime}$ with respect to a given similarity threshold $s\in[0,1]$. OV focuses only on the equivalence relation; subsumption mappings are beyond the scope of versioning.
@@ 3.2 OM4OV Pipeline
Given ontologies $O$ and $O^{\prime}$ and similarity threshold $s\in[0,1]$, the OV task is to find $A_{match}$ and $A_{non-match}$ as follows:
$A_{match}$ $=\{(e_{1},e_{2},c)\mid e_{1}\in O,e_{2}\in O^{\prime},s\leq c\leq 1\ \mbox{and there is no }(e_{1},e_{2},c^{\prime})\mbox{ with }c^{\prime}>c\}$ (1)
$A_{non-match}$ $=\{e\mid e\in O\cup O^{\prime}\mbox{and there is no }(e_{1},e_{2},c)\in A_{match}\mbox{ with }e\in\{e_{1},e_{2}\}\}$
For convenience, we will sometimes refer to the elements of $A_{match}$ as entities rather than alignments, by which we mean entities that are mentioned in mappings in $A_{match}$.
The remain and update entities are entities successfully matched over different ontology versions. The remain entities can be considered to be the exact (trivial) equivalent string matches between different ontology versions. On the other hand, if the convention used to represent an entity name changes (e.g. from EyelidTarsus to eyelid_tarsus), we would expect an OV system to classify the change as an update by modelling the confidence $c$ as a value greater than $s$ but less than 1. The update entities are, in general, the closest (non-trivial) near-equivalent matches between different ontology versions.
@@
$A_{remain}(A\odot)$ and $A_{update}(A\otimes)$ are defined as:
$A\odot=\{(e_{1},e_{2},c)\mid(e_{1},e_{2},c)\in A_{match}\mbox{ and }c=1\}$ (2)
$A\otimes=\{(e_{1},e_{2},c)\mid(e_{1},e_{2},c)\in A_{match}\mbox{ and }c<1\}$
We can see that $A_{match}=A\odot\cup A\otimes$
The add and delete entities are actually non-matched entities between different ontology versions. The add entities are the entities in $O^{\prime}$ which have no matches in $O$. Similarly, we can interpret the delete entities as entities in $O$ which have no matches in $O^{\prime}$.
@@
$A_{add}(A\oplus)$ and $A_{delete}(A\ominus)$ are defined as:
$A\oplus=\{e\mid e\in O^{\prime}\mbox{ and }e\in A_{non-match}\}$ (3)
We can see that $A_{non-match}=A\oplus\cup A\ominus$. It is expected that $A\oplus$ and $A\ominus$ are disjoint, but this would be violated when a faulty OM4OV pipeline determines that an equivalent pair of entities in each ontology does not match.
@@ 4.1 Dataset Construction
There is a dearth of benchmark datasets for evaluating OV. The Ontology Alignment Evaluation Initiative (OAEI) contains several datasets related to OM tasks, but none are specifically designed for OV tasks. We propose an approach to constructing synthetic OV datasets from real-world OM datasets. Figure 2 illustrates the generation of OM4OV datasets.
Figure 2: Generation of OM4OV datasets from initial OM inputs of source ontology, target ontology, and reference alignment, and output versioning ontologies and reference alignments with four categories.
@@
The original OAEI datasets for OM provide two ontologies, the source $O_{s}$ and the target $O_{t}$. We choose either one as the ontology for versioning and duplicate it into $O$ and $O^{\prime}$. Another ontology will be used later to provide references for synthetic update entities. We duplicate this ontology into $O_{r}$.
There are four possible entity changes in OV tasks: remain, update, add, and delete. We retrieve all ontology entities from $O$ or $O^{\prime}$ (since they are identical at this stage). Each ontology entity will be randomly assigned to one of these categories, and the four sets are pairwise disjoint by construction. In real-world OV, entity updates are relatively infrequent. We allow a user-defined parameter to fix the proportion of update entities in the synthetic dataset, defaulting to 25%. The update entities are randomly selected from the replacement corpus, so the number of update entities is limited by the size of the replacement corpus, which may be less than the user parameter value. For this reason, selection of update entities is prioritised over other categories, and excess entities are redistributed evenly over the other three categories.
For each ontology entity $e$ assigned to a category, the corresponding dataset operations are described as follows. While there is no operation for remain entities, entities assigned to add and delete are deleted from all triples that mention them in $O$ or $O^{\prime}$ respectively. Each update mapping requires that all occurrences of its domain entity in triples in $O^{\prime}$ are replaced by the corresponding co-domain entity of the mapping. We generate four corresponding versioning references based on the entity assignments. These files provide the ground truth for evaluation.
To ensure the replaced entity is reasonable and reflects real-world facts, we use equivalent classes or properties specified in reference.xml. For example, we could replace cmt:ProgramCommitteeChair with its equivalent class confof:Chair_PC. This operation is cascaded, meaning that we update not only the entity name but also its related triples, such as changing the semantic information from (cmt:ProgramCommitteeChair, rdfs:subClassOf, cmt:ProgramCommitteeMember) to (confof:Chair_PC, rdfs:subClassOf, confof:Person). While this procedure can be unlimited, we restrict it to extract only the triples directly related to the entity (i.e. a 1-hop subgraph) in the ontology $O_{r}$. For entities whose names are unique identifiers or codes (and not textually meaningful names), we retrieve their meaningful descriptions from annotation properties. For example, we extract an additional annotation triple (mouse:MA_0000270, rdfs:label, “eyelid tarsus”) for the entity mouse:MA_0000270. After extraction, we also need to update the URI from “confof” to “cmt” where the new entity URI will be cmt:Chair_PC and the replacement triple will be (cmt:Chair_PC, rdfs:subClassOf, cmt:Person). Finally, we add these new triples to the existing ontology $O^{\prime}$. Table 1 summarises the detailed steps of the dataset operation for update entities.
@@
Unlike the original OAEI datasets used for OM, we introduce randomness to ensure that the synthetic OAEI datasets for OV are different each time they are constructed. This suits the dynamic nature of OV, where the changes vary between different versions. For this reason, we consider the new OAEI datasets for OV more like a testbed, as they can simulate a variety of situations for OV tasks. When reproducibility is required, we use a fixed seed to control the randomness. Table 2 details the OAEI tracks selected for the OV testbed. The current version of the OV testbed contains a total of 11 distinct ontologies from three different OAEI tracks. The anatomy track contains two large ontologies, while the MSE track has two medium ontologies, and the conference track has 7 small ontologies. We exclude EMMO [9] from the MSE track because it assigns opaque symbolic codes to entity names, using a naming convention that differs from MaterialInformation [3] and MatOnto [18], which use meaningful entity names. Figure 3 shows the number of entities in each ontology in each track.
@@
Track Domain # of Ontologies Anatomy Human and Mouse Anatomy 2 Conference Research Conference 7 MSE Materials Science and Engineering 2
@@
4.2 System Demonstration
We choose the matching system Agent-OM and its extended version Agent-OV to demonstrate the OM4OV pipeline. Agent-OM is one of the flagship OM systems powered by large language models (LLMs) and multi-agent architectures. Its foundation framework is designed for traditional OM tasks. We extend Agent-OM with the OM4OV pipeline so that it can be used for OV tasks. Inherited from Agent-OM, Agent-OV supports a wide range of LLMs, including commercial API-accessed LLMs OpenAI GPT [28] and Anthropic Claude [2], as well as open-source LLMs Meta Llama [24], Alibaba Qwen [1], Google Gemma [11], and ChatGLM [36]. Similarly to Agent-OM, similarity_thresholdsimilarity\_threshold and $top@k$ are the adjustable key hyperparameters for Agent-OV.